# High-inclination Centaur reservoirs beyond Neptune

### v1

@@ 6 Injected Centaur orbits
The analysis of orbit injection in the three-body problem indicates that a Centaur orbit acquires its largest inclination at its minimum semi-major axis as it follows its Tisserand inclination pathway inside the planet’s orbit (Paper III). This makes the inclination at minimum semi-major axis a useful diagnostic tool to access pre-injection original inclinations far from the planet (3) through the Tisserand parameter.
@@
Figure 8 shows the Centaurs’ inclination at minimum semi-major axis as a function of the Tisserand parameter for hot TNO reservoirs and the different initial semi-major axes. The statistics for an initial semi-major axis include all initial perihelia. The top panel displays the mean inclinations and the bottom panel the standard deviations. The dashed curves are those of $a=40$ au (magenta), $a=200$ au (brown) and $a=500$ au (black) that lie outside the stable region. The remaining semi-major axes are not identified in the figure because the curves are close but generally from top to bottom the curves represent increasing semi-major axes. The color codes are those of Figures 3 and 5.
@@
Figure 8: Centaur inclination at minimum semi-major axis of initially hot TNOs as a function of the Tisserand parameter $T$ (and $I_{\infty}$) for semi-major axes $50\leq a$ (au) $\leq 140$ (solid curves) and $a=40$ au (dashed magenta curve), $a=200$ au (dashed black curve) and $a=500$ au (dashed brown curve). The average inclination, $\langle I\rangle$, and standard deviation, $\sigma_{I}$, are shown in the top and bottom panels respectively.
@@
Figure 9: Semi-major axis-averaged Centaur inclination at minimum semi-major axis of cold TNOs (blue) and hot TNOs (red) as a function of the Tisserand parameter $T$ (and $I_{\infty}$). Dashed blue and red lines are the corresponding standard deviations. The solid black lines are $\langle I\rangle=110^{\circ}-30^{\circ}T$ for $T\geq-0.67$ and $\langle I\rangle=135^{\circ}$ otherwise.
@@
Figure 10: Centaur minimum semi-major axis of initially hot TNOs as a function of the Tisserand parameter $T$ (and $I_{\infty}$) for semi-major axes $50\leq a$ (au) $\leq 140$ (solid lines) and $a=40$ au (dashed magenta curve), $a=200$ au (dashed brown line) and $a=500$ au (dashed black line). The average semi-major axis, $\langle a_{\rm min}\rangle$, and standard deviation, $\sigma_{a}$, are shown in the top and bottom panels respectively.
@@
Figure 11: Average Centaur minimum semi-major axis of cold TNOs (blue) and hot TNOs (red) as a function of the Tisserand parameter $T$ (and $I_{\infty}$). Dashed blue and red curves are the corresponding standard deviations. The solid black line is $\langle a_{\rm min}\rangle$(au) =18.5T=18.5-T. The solid orange curve is the three-body reflection semi-major axis of Neptune (4). The planets’ positions are indicated with dashed black lines.
@@
The mean inclinations remarkably share the same functional dependence on the Tisserand parameter regardless of semi-major axis. There is a Tisserand value around $T\sim-0.5$, for each semi-major axis, below which the mean inclination is constant and beyond which inclination decreases linearly with the Tisserand parameter. The constant inclination value for T<0.5T<-0.5 decreases form $\sim 150^{\circ}$ for $40$ au to $\sim 120^{\circ}$ for $140$ au.
@@
Figure 9 shows the semi-major axis-averaged inclinations at minimum semi-major axis of Centaurs originating in the stable region as a function of the Tisserand parameter for hot and cold TNO reservoirs along with their standard deviations. The initially hot TNO Centaur-inclination is approximated by $\langle I\rangle=110^{\circ}-30^{\circ}T$ for $T\geq-0.67$ and $\langle I\rangle=135^{\circ}$ otherwise. The cold TNO Centaur-inclination has a similar dependence. The inclination standard deviation increases steadily from $\sigma_{I}\sim 25^{\circ}$ at $T=2.8$ to $\sigma_{I}\sim 45^{\circ}$ for T=2T=-2.
@@
The Centaur inclination at minimum semi-major axis shows that high-inclination Centaurs may not be produced by reservoirs with arbitrary initial inclinations. For instance, a low inclination reservoir with $T=2.8$ has $\langle I\rangle\sim 30^{\circ}$. Retrograde Centaurs with inclinations $100^{\circ}$ and $155^{\circ}$ respectively lie at $3\sigma_{I}$ and $5\sigma_{I}$ from the mean value. This explains the inability of relaxation models of the early planar protoplanetary disk to account for the numbers of high-inclination Centaurs. On the other hand, reservoirs in the polar corridor may produce polar and retrograde inclination Centaurs with statistical significance.
@@
The minimum semi-major axes (mean and standard deviation) of hot TNOs, that achieved a Centaur orbit are shown in Figure 10 as a function of the Tisserand parameter for different initial semi-major axes. The minimum semi-major axis is averaged over all initial perihelia. The semi-major axes are not indicated because the curves are close. Centaur injection occurs across all values of the Tisserand parameter in contrast to the dynamics of the three-body problem. The minimum semi-major axes decrease with increasing $T$ in a similar fashion. The large variations of the TNO ensembles with an initial $a=500$ au comes for the small statistics of injected Centaurs. The minimum semi-major axes averaged over the different initial semi-major axes of Centaurs originating in the stable region are shown in Figure 11 for hot and cold TNOs. The two curves are similar and may be approximated by the linear expression $\langle a_{\rm min}\rangle$(au)=18.5T=18.5-T with a mostly constant standard deviation $\sigma_{a}\sim 8$ au. The three-body reflection semi-major axis of Neptune (4) shown in Figure 11 is of the same order as $\langle a_{\rm min}\rangle$ for $T>0.4$ only as it does not account for the presence of the giant planets.
@@
Figure 12: Centaur perihelion at minimum semi-major axis of initially hot TNOs as a function of the Tisserand parameter $T$ (and $I_{\infty}$) for semi-major axes $50\leq a$ (au) $\leq 140$ (solid lines) and $a=40$ au (dashed magenta curve), $a=200$ au (dashed brown line) and $a=500$ au (dashed black line). The average semi-major axis, $\langle q\rangle$, and standard deviation, $\sigma_{q}$, are shown in the top and bottom panels respectively.
@@
Figure 13: Average Centaur perihelion at minimum semi-major axis of cold TNOs (blue) and hot TNOs (red) as a function of the Tisserand parameter $T$ (and $I_{\infty}$). Dashed blue and red curves are the corresponding standard deviations. The solid black line is $\langle q\rangle$(au) =6.9+0.4T=6.9+0.4T. The planets’ positions are indicated with dashed black lines.
@@
The perihelia at minimum semi-major axis (mean and standard deviation) of hot TNOs that achieved a Centaur orbit are shown in Figure 12 as a function of the Tisserand parameter for different initial semi-major axes. Mean perihelia and their standard deviations increase with increasing Tisserand parameter (decreasing initial inclination). It is also found that perihelia decrease with increasing initial semi-major axis. In particular, hot TNOs with $a=500$ au are injected as Centaurs with perihelia around Jupiter’s orbit. The semi-major axis averaged Centaur perihelia originating in the stable region are shown in Figure 13. The hot TNO’s mean perihelion as a function of the Tisserand parameter follows the linear relation $\langle q\rangle$(au) =6.9+0.4T=6.9+0.4T. The standard deviation increases from 3 au at T=2T=-2 to 4 au at $T=2.8$.
@@ 7 Conclusion
We set out to confirm the possible existence of current TNO reservoirs beyond Neptune’s orbit that supply high-inclination Centaurs to the inner Solar System. This confirmation is based on simulating the gravitational perturbations of TNO reservoirs by the four giant planets, passing stars and the Galactic tide. The initial extent of the TNO reservoirs made use of two previous results. The first is the existence of a stable region in the semi-major range axis [40:160] au in the Sun-Neptune-TNO three-body problem (Paper IV). The second result is the proximity of past Centaur perihelia to Neptune’s orbit found in the 4.5 Gyr time-backward simulations of 19 real high-inclination Centaurs using 20 million clones (Papers I and II). These extents were enlarged to better understand Centaur injection. The perihelion range was enlarged to [32:50] au, and a cold population with circular orbits was added to examine if Centaur injection is strongly dependent on perihelion. The Tisserand range was enlarged to [2-2:2.8] or equivalently to TNO inclinations far from Neptune in the range $8^{\circ}$$135^{\circ}]$ to examine Centaur injection beyond the polar corridor’s retrograde inclination boundary in the Sun-Neptune-TNO three-body problem, $I_{\infty}\sim 110^{\circ}$ (T=1T=-1), and also at low inclinations $I_{\infty}\sim 8^{\circ}$ ($T=2.8$) of relevance to the relaxation of the early protoplanetary disk.
@@
The numerical simulations confirmed the existence of current TNO reservoirs that provide high-inclination Centaurs to the Solar System as $\sim 40$% of the initial hot TNOs in the polar corridor survive 4.5 Gyr (Figure 4) in the semi-major axis range [50:140] au. It was also found that two specific Tisserand parameter values $T=0.5$ and T=1.5T=-1.5 (initial inclinations at infinity $80^{\circ}$ and $122^{\circ}$) have the largest current populations (Figure 3). A finer analysis unveiled regions in perihelion space where the dynamical time is 4.5 Gyr and hence surviving hot TNOs constitute more than half the original reservoir (Figure 6).
@@
The role of the Tisserand inclination pathway with respect to Neptune in the reservoirs’ evolution is confirmed (Figure 2). The presence of all the giant planets modifies the polar corridor’s structure as Saturn becomes the main perturber of a fraction of TNO material adding secondary Tisserand inclination pathways to the polar corridor (Figure 2). These pathways are defined only by Saturn regardless of the TNOs’ initial Tisserand parameters. Neptune’s reflection ability of incoming TNOs is also modified by the the giant planets as Centaur injection occurs outside the polar corridor’s high-inclination boundary of T=1T=-1. Although the presence of more planets may cause significant departures from three-body predictions, the Tisserand inclination pathway concept is robust as TNO dynamics remain based on it.
@@
The existence of linear relations with respect to the Tisserand parameter that describe, the number of injected Centaurs (Figure 5), the Centaur inclination at minimum semi-major axis (Figure 9) and the Centaur minimum semi-major axis (Figure 11), affirms the role of the Tisserand parameter as the main control parameter in TNO dynamics. These relations were obtained by averaging over initial perihelia and semi-major axes for hot and cold TNOs. That the two populations follow similar Tisserand parameter-based linear relations indicates that the Centaur injection process is largely independent of the initial perihelia. The dependence of the mean Centaur perihelia on the initial semi-major axis showed the larger the TNO initial semi-major axis, the smaller the perihelion (Figure 12). Beyond this characterization of the injected Centaur orbits of high-inclination-TNO reservoirs, the linear relations will help discriminate between different high-inclination Centaur origins such as the Oort cloud component formed by the protoplanetary disk and the excitation caused by additional planets beyond Neptune provided that Centaur injection is analyzed in terms of the Tisserand parameter using orbital elements at minimum semi-major axis.
@@
The most compelling result regarding Centaur injection is perhaps how the Centaur inclination at minimum semi-major axis depends on the Tisserand parameter and how it is independent of the initial semi-major axis (Figure 8). It explains why high-inclination Centaurs may not be produced with statistical significance from a low inclination reservoir such as the early protoplanetary disk, and why matter in the polar corridor may yield Centaurs with retrograde orbits.
@@
Matter in the stable region at 4.5-4.5 Gyr is unlikely to originate in the early planetesimal disk because it was flat and did not extend beyond $\sim 30$ au in order to explain Neptune’s current location (Gomes04). The natural source for that matter is the stellar environment of the early Solar System. Simulations of interstellar matter capture in the Sun’s birth cluster form the inner Oort cloud with an inner edge near 100 au (Brasser12b) which falls in the middle of the stable region found in this work. Simulations of star encounters in the presence of protoplanetary disks show that it is possible to capture planetesimals at high inclination while preserving a star’s flat planetesimal disk (Hands19). After matter exchange in the Solar System’s early environment has settled, the stable region of this work is where the TNO reservoirs may currently be found nearest to the giant planets. In similar way to the Trojans asteroids found at Jupiter’s stable Lagrange points that are relics of early planet formation, the high-inclination reservoirs are relics of the Solar System’s interactions with the stars of its birth cluster. The Legacy Survey of Space and Time (LSST) of the Vera Rubin Observatory will be able to constrain their extent and population size through its coverage of high ecliptic latitudes necessary to detect eccentric polar TNOs that spend more time above the ecliptic plane (Kurlander25).

### v2

@@ 6 Injected Centaur orbits
The analysis of orbit injection in the three-body problem indicates that a Centaur orbit acquires its largest inclination at its minimum semimajor axis as it follows its Tisserand inclination pathway inside the planet’s orbit (Paper III). This makes the inclination at minimum semimajor axis a useful diagnostic tool for accessing pre-injection original inclinations far from the planet (3) through the Tisserand parameter.
@@
Figure 8 shows the Centaurs’ inclination at minimum semimajor axis as a function of the Tisserand parameter for hot TNO reservoirs and the different initial semimajor axes. The statistics for an initial semimajor axis include all initial perihelia. The top panel displays the mean inclinations and the bottom panel the standard deviations. The different initial semimajor axes are not indicated because the curves are close.
@@
Figure 8: Centaur inclination at minimum semimajor axis of initially hot TNOs as a function of $T$ (and $I_{\infty}$) for semimajor axes $50\leq a$ (au) $\leq 140$ (solid curves) inside the stable region, and a=40, 200, 500a=40,\ 200,\ 500 au (dashed curves) outside the stable region. The color codes are those of Figs. 3 and 5. The average inclination, $\langle I\rangle$, and standard deviation, $\sigma_{I}$, are shown in the top and bottom panels, respectively.
@@
Figure 9: Semimajor axis-averaged Centaur inclination at minimum semimajor axis of cold TNOs (blue) and hot TNOs (red) as a function of $T$ (and $I_{\infty}$). Dashed blue and red lines are the corresponding standard deviations. The solid black lines are $\langle I\rangle=110^{\circ}-30^{\circ}T$ for $T\geq-0.67$ and $\langle I\rangle=135^{\circ}$ otherwise.
@@
Figure 10: Centaur minimum semimajor axis of initially hot TNOs as a function of $T$ (and $I_{\infty}$) for semimajor axes $50\leq a$ (au) $\leq 140$ (solid curves) inside the stable region, and a=40, 200, 500a=40,\ 200,\ 500 au (dashed curves) outside the stable region. The color codes are those of Figs. 3 and 5. The average semimajor axis, $\langle a_{\rm min}\rangle$, and standard deviation, $\sigma_{a}$, are shown in the top and bottom panels, respectively.
@@
Figure 11: Average Centaur minimum semimajor axis of cold TNOs (blue) and hot TNOs (red) as a function of $T$ (and $I_{\infty}$). Dashed blue and red curves are the corresponding standard deviations. The solid black line is $\langle a_{\rm min}\rangle$(au) =18.5T=18.5-T. The solid orange curve is the three-body reflection semimajor axis of Neptune (4). The planets’ positions are indicated with dashed black lines.
@@
The mean inclinations remarkably share the same functional dependence on the Tisserand parameter regardless of semimajor axis. There is a Tisserand value around $T\sim-0.5$, for each semimajor axis, below which the mean inclination is constant and beyond which inclination decreases linearly with the Tisserand parameter. The constant inclination value for T<0.5T<-0.5 decreases form $\sim 150^{\circ}$ for $40$ au to $\sim 120^{\circ}$ for $140$ au.
@@
Figure 9 shows the semimajor axis-averaged inclinations at minimum semimajor axis of Centaurs originating in the stable region as a function of the Tisserand parameter for hot and cold TNO reservoirs along with their standard deviations. The initially hot TNO Centaur-inclination is approximated by $\langle I\rangle=110^{\circ}-30^{\circ}T$ for $T\geq-0.67$ and $\langle I\rangle=135^{\circ}$ otherwise. The cold TNO Centaur-inclination has a similar dependence. The inclination standard deviation increases steadily from $\sigma_{I}\sim 25^{\circ}$ at $T=2.8$ to $\sigma_{I}\sim 45^{\circ}$ for T=2T=-2.
@@
The Centaur inclination at minimum semimajor axis shows that high-inclination Centaurs may not be produced by reservoirs with arbitrary initial inclinations. For instance, a low-inclination reservoir with $T=2.8$ has $\langle I\rangle\sim 30^{\circ}$. Retrograde Centaurs with inclinations $100^{\circ}$ and $155^{\circ}$ respectively lie at $3\sigma_{I}$ and $5\sigma_{I}$ from the mean value. This explains the inability of relaxation models of the early planar protoplanetary disk to account for the numbers of high-inclination Centaurs. On the other hand, reservoirs in the polar corridor can produce polar and retrograde inclination Centaurs with statistical significance.
@@
The minimum semimajor axes (mean and standard deviation) of hot TNOs that achieved a Centaur orbit are shown in Fig. 10 as a function of the Tisserand parameter for different initial semimajor axes. The minimum semimajor axis is averaged over all initial perihelia. Centaur injection occurs across all values of the Tisserand parameter in contrast to the dynamics of the three-body problem. The minimum semimajor axes decrease with increasing $T$ in a similar fashion. The large variations of the TNO ensembles with an initial $a=500$ au comes for the small statistics of injected Centaurs. The minimum semimajor axes averaged over the different initial semimajor axes of Centaurs originating in the stable region are shown in Fig. 11 for hot and cold TNOs. The two curves are similar and can be approximated by the linear expression $\langle a_{\rm min}\rangle$(au)=18.5T=18.5-T with a mostly constant standard deviation $\sigma_{a}\sim 8$ au. The three-body reflection semimajor axis of Neptune (4) shown in Fig. 11 is of the same order as $\langle a_{\rm min}\rangle$ only for $T>0.4$ as it does not account for the presence of the giant planets.
@@
Figure 12: Centaur perihelion at minimum semimajor axis of initially hot TNOs as a function of $T$ (and $I_{\infty}$) for semimajor axes $50\leq a$ (au) $\leq 140$ (solid curves) inside the stable region, and a=40, 200, 500a=40,\ 200,\ 500 au (dashed curves) outside the stable region. The color codes are those of Figs. 3 and 5. The average semimajor axis, $\langle q\rangle$, and standard deviation, $\sigma_{q}$, are shown in the top and bottom panels, respectively.
@@
Figure 13: Average Centaur perihelion at minimum semimajor axis of cold TNOs (blue) and hot TNOs (red) as a function of $T$ (and $I_{\infty}$). Dashed blue and red curves are the corresponding standard deviations. The solid black line is $\langle q\rangle$(au) =6.9+0.4T=6.9+0.4T. The planets’ positions are indicated with dashed black lines.
@@
The perihelia at minimum semimajor axis (mean and standard deviation) of hot TNOs that achieved a Centaur orbit are shown in Fig. 12 as a function of the Tisserand parameter for different initial semimajor axes. Mean perihelia and their standard deviations increase with increasing Tisserand parameter (decreasing initial inclination). It is also found that perihelia decrease with increasing initial semimajor axis. In particular, hot TNOs with $a=500$ au are injected as Centaurs with perihelia around Jupiter’s orbit. The semimajor axis averaged Centaur perihelia originating in the stable region are shown in Fig. 13. The hot TNO’s mean perihelion as a function of the Tisserand parameter follows the linear relation $\langle q\rangle$(au) =6.9+0.4T=6.9+0.4T. The standard deviation increases from 3 au at T=2T=-2 to 4 au at $T=2.8$.
@@ 7 Conclusion
We set out to confirm the possible existence of current TNO reservoirs beyond Neptune’s orbit that supply high-inclination Centaurs to the inner Solar System. We did so by simulating the gravitational perturbations of TNO reservoirs by the four giant planets, passing stars, and the Galactic tide. We based the initial extent of the TNO reservoirs on two previous results. The first is the existence of a stable region in the semimajor range axis [40:160] au in the Sun-Neptune-TNO three-body problem (Paper IV). The second is the proximity of past Centaur perihelia to Neptune’s orbit found in the 4.5 Gyr time-backward simulations of 19 real high-inclination Centaurs that involved 20 million clones (Papers I and II). These extents were enlarged to better understand Centaur injection: the perihelion range was enlarged to [32:50] au, and a cold population with circular orbits was added to examine if Centaur injection is strongly dependent on perihelion. The Tisserand range was enlarged to [2-2:2.8], or equivalently to TNO inclinations far from Neptune in the range $8^{\circ}$$135^{\circ}],$ to examine Centaur injection beyond the polar corridor’s retrograde inclination boundary in the Sun-Neptune-TNO three-body problem, $I_{\infty}\sim 110^{\circ}$ (T=1T=-1), and also at low inclinations of relevance to the relaxation of the early protoplanetary disk, $I_{\infty}\sim 8^{\circ}$ ($T=2.8$).
@@
The numerical simulations confirmed the existence of current TNO reservoirs that provide high-inclination Centaurs to the Solar System with $\sim 40$% of the initial hot TNOs in the polar corridor surviving 4.5 Gyr (Fig. 4) in the semimajor axis range [50:140] au. It was also found that two specific Tisserand parameter values, $T=0.5$ and T=1.5T=-1.5 (initial inclinations at infinity of $80^{\circ}$ and $122^{\circ}$), have the largest current populations (Fig. 3). A finer analysis unveiled regions in perihelion space where the dynamical time is 4.5 Gyr and, hence, surviving hot TNOs constitute more than half the original reservoir (Fig. 6).
@@
The role of the Tisserand inclination pathway with respect to Neptune in the reservoirs’ evolution is confirmed (Fig. 2). The presence of all the giant planets modifies the polar corridor’s structure as Saturn becomes the main perturber of a fraction of TNO material, adding secondary Tisserand inclination pathways to the polar corridor (Fig. 2). These pathways are defined only by Saturn regardless of the TNOs’ initial Tisserand parameters. Neptune’s reflection ability of incoming TNOs is also modified by the giant planets because Centaur injection occurs outside the polar corridor’s high-inclination boundary of T=1T=-1. Although the presence of more planets causes some significant departures from three-body predictions, the Tisserand inclination pathway concept is robust as TNO dynamics remain based on it.
@@
The existence of linear relations with respect to the Tisserand parameter that describe the number of injected Centaurs (Fig. 5), the Centaur inclination at minimum semimajor axis (Fig. 9), and the Centaur minimum semimajor axis (Fig. 11) affirms the role of the Tisserand parameter as the main control parameter in TNO dynamics. These relations were obtained by averaging the orbital elements over initial perihelia and semimajor axes for hot and cold TNOs. That the two populations follow similar Tisserand parameter-based linear relations indicates that the Centaur injection process is largely independent of the initial perihelia. The dependence of the mean Centaur perihelia on the initial semimajor axis shows that the larger the TNO initial semimajor axis, the smaller the perihelion (Fig. 12). Beyond this characterization of the injected Centaur orbits of high-inclination-TNO reservoirs, the linear relations will help us discriminate between different high-inclination Centaur origins, such as the Oort cloud component formed by the protoplanetary disk and the excitation caused by additional planets beyond Neptune — provided that Centaur injection is analyzed in terms of the Tisserand parameter using orbital elements at minimum semimajor axis.
@@
The most compelling result regarding Centaur injection is perhaps how the Centaur inclination at minimum semimajor axis depends on the Tisserand parameter and how it is independent of the initial semimajor axis (Fig. 8). This explains why high-inclination Centaurs are not produced with statistical significance from a low-inclination reservoir such as the early protoplanetary disk, and why matter in the polar corridor can yield Centaurs with retrograde orbits.
@@
Matter in the stable region at 4.5-4.5 Gyr is unlikely to originate in the early planetesimal disk because it was flat, and to explain Neptune’s current location, it must not have extended beyond $\sim 30$ au (Gomes04). The natural source of that matter is the stellar environment of the early Solar System. According to simulations of interstellar matter capture in the Sun’s birth cluster, the inner Oort cloud formed with an inner edge near 100 au (Brasser12b), which falls in the middle of the stable region found in this work. Simulations of star encounters in the presence of protoplanetary disks show that it is possible to capture planetesimals at high inclinations while preserving a star’s flat planetesimal disk (Hands19). After matter exchange in the Solar System’s early environment has settled, the stable region of this work is where the TNO reservoirs can currently be found nearest to the giant planets. Like the Trojans asteroids found at Jupiter’s stable Lagrange points that are relics of early planet formation, the high-inclination reservoirs are relics of the Solar System’s interactions with the stars of its birth cluster. The Legacy Survey of Space and Time (LSST) of the Vera Rubin Observatory will be able to constrain their extent and population size through its coverage of high ecliptic latitudes necessary to detect eccentric polar TNOs that spend more time above the ecliptic plane (Kurlander25).